# Scalar Field Dominated Cosmology with Woods-Saxon Like Potential

### v1

@@ 4 Cosmographic Parameters
Figure 2: The evolution of Hubble parameter($H$) against redshift($z$) for the WSQ model and $\Lambda$CDM using the best-fit parameters for SNIa+OHD dataset.
@@
The evolution of the equation of state parameter $\omega$ of the scalar field $\phi$ is dynamical, while it is a constant in the $\Lambda$CDM with a value $=-1$ during the entire evolution of the Universe. The evolution of $\omega$ is presented in Fig. 3. It is evident that for $z>4.43$, the scalar field together with matter contributed to the deceleration. For $z<4.43$, $-\frac{1}{3}<\omega<-1$, showing that the scalar field started contributing to the accelerated expansion, representing a freezing model of quintessence. The present value of $\omega$ is obtained as $-0.9617$.
@@
Figure 3: Evolution of the equation of state ($\omega$) against redshift($z$), plotted with the best-fit parameters obtained for SNIa+OHD dataset.
Figure 4: The age of the Universe against redshift($z$) is plotted for the WSQ model and $\Lambda$CDM using the best-fit parameters obtained for the SNIa + OHD dataset.
@@
Figure 5: Evolution of matter and dark energy density in the WSQ model plotted with best-fit parameters obtained for the SNIa + OHD dataset.
@@
$q=-1-\frac{\dot{H}}{H^{2}}.$ (25)
In terms of redshift, we can express Eq. (25) as,
$q=-1+\frac{(1+z)}{H}\frac{dH}{dz}.$ (26)
Figure 6: Evolution of deceleration parameter $q$ with redshift $z$ for the WSQ model and $\Lambda$CDM, plotted with best-fit parameters obtained for the SNIa+OHD dataset.
A negative value for $q$ would signify accelerated expansion of the Universe. In the WSQ model, we obtain the present deceleration parameter value of $q_{0}=-0.501$, supporting the current accelerated expansion of the Universe. The redshift at which the Universe transitioned from decelerating to accelerating expansion is estimated to be $z_{T}=0.662$ as compared to a value of $0.70$ obtained from the $\Lambda$CDM. The WSQ model is successful in explaining the observed accelerated expansion of the Universe and the transition from decelerating to accelerating phase. The evolution of deceleration parameter with redshift is plotted in Fig. 6.
@@
$r=\frac{1}{a}\frac{d^{3}a}{dt^{3}}\left[\frac{1}{a}\frac{da}{dt}\right]^{-3}=\frac{\dddot{a}}{aH^{3}}.$ (27)
Figure 7: The statefinder trajectory is plotted using the best-fit parameters from the SNIa + OHD dataset. The fixed point of $\Lambda$CDM is also marked.
We can represent $r$ in terms of the dimensionless Hubble parameter $h$ while introducing a new variable $N=log_{e}(a)$, we obtain,
$r=\frac{1}{2h^{2}}\frac{d^{2}h^{2}}{dN^{2}}+\frac{3}{2h^{2}}\frac{dh^{2}}{dN}+1.$ (28)
@@
$s=\frac{r-1}{3(q-1/2)}.$ (29)
On expressing $q$ presented in Eq. (25) in terms of $N$ and substituting in Eq. (29), we obtain
$s=-\frac{\frac{1}{2h^{2}}\frac{d^{2}h^{2}}{dN^{2}}+\frac{3}{2h^{2}}\frac{dh^{2}}{dN}}{\frac{3}{2h^{2}}\frac{dh^{2}}{dN}+\frac{9}{2}}.$ (30)
The statefinder trajectory distinguishes the behaviour of different types of dark energy models. The trajectory is plotted in Fig. 7.
@@
The statefinder trajectory reveals the quintessence behaviour in the present and future while the field behaved as a Chaplygin gas in the past. The present value of the statefinder diagnostic pair is obtained as ($r,s$) = ($0.8309,0.0562$), different from the $\Lambda$CDM fixed point ($1,0$).
@@ 5 Conclusion
In the presented work, we examined the proposed Woods Saxon quintessence model with observational constraints. We have applied MCMC methods for the Bayesian inference of the model parameters using the data combination SNIa + OHD. Our analysis shows that the model is consistent with the observational data and the obtained values of parameters are also consistent with that obtained with $\Lambda$CDM for the same dataset. The obtained best-fit values of model parameters are $H_{0}=68.10\pm 0.74$, $\Omega_{m_{0}}=0.31\pm 0.01$ and $y_{0}=0.83\pm 0.01$, which is in close agreement with $\Lambda$CDM predictions. The WSQ model is preferred over the $\Lambda$CDM according to $\chi^{2}_{min}$ and AIC, while according to the BIC criteria, the $\Lambda$CDM is slightly preferred over the WSQ model, because of the extra parameter.
The study of evolution of different cosmological parameters also revealed the consistency of the WSQ model in explaining the late-time background evolution of the Universe. The evolution of $\omega$ shows a freezing model of quintessence. Age of the Universe is estimated to be $13.6$ Gyr, which is slightly less than that predicted by the $\Lambda$CDM. The present value of the deceleration parameter obtained with the WSQ model is negative, showing that the model predicts present day accelerated expansion. The Universe transitioned from decelerating to accelerating is at a redshift $z_{T}=0.662$. The statefinder trajectory reveals that the field behaves as a Chaplygin gas before reaching the $\Lambda$CDM fixed point and as a quintessence after this point.
In summary, the model is consistent with the observations to good extent and can be considered as a potential alternative to the $\Lambda$CDM to explain the late phase acceleration of the Universe.

### v2

@@ 4 Cosmographic Parameters
Figure 2: The evolution of Hubble parameter($H$) against redshift($z$) for the WSQ model and $\Lambda$CDM using the best-fit parameters for SNIa+OHD datasets.
The evolution of the equation of state parameter $\omega$ of the scalar field $\phi$ is dynamical, while it is a constant in the $\Lambda$CDM with a value $=-1$ during the entire evolution of the Universe. The evolution of $\omega$ is presented in Fig. 3. It is evident that for $z>4.43$, the scalar field together with matter contributed to the deceleration. For $z<4.43$, the equation of state $-\frac{1}{3}<\omega<-1$, which shows the scalar field started contributing to the accelerated expansion, representing a freezing model of quintessence. The present value of $\omega$ is obtained as $-0.9617$. Furthermore, $\omega$ attains a value equal to 1 at larger redshifts, showing a stiff fluid nature of the scalar field at higher redshifts.
Figure 3: Evolution of the equation of state ($\omega$) against redshift($z$), plotted with the best-fit parameters obtained for SNIa+OHD datasets.
Figure 4: The age of the Universe against redshift($z$) is plotted for the WSQ model and $\Lambda$CDM using the best-fit parameters obtained for the SNIa + OHD datasets.
@@
Figure 5: Evolution of matter and dark energy density in the WSQ model plotted with best-fit parameters obtained for the SNIa + OHD datasets.
@@
$q=-1-\frac{\dot{H}}{H^{2}}.$ (4.2)
In terms of redshift, we can express Eq. (4.2) as,
$q=-1+\frac{(1+z)}{H}\frac{dH}{dz}.$ (4.3)
A negative value for $q$ would signify accelerated expansion of the Universe. In the WSQ model, we obtain the present deceleration parameter value of $q_{0}=-0.501$, supporting the current accelerated expansion of the Universe. The redshift at which the Universe transitioned from decelerating to accelerating expansion is estimated to be $z_{T}=0.662$. The WSQ model is successful in explaining the observed accelerated expansion of the Universe and the transition from decelerating to accelerating phase. The evolution of deceleration parameter with redshift is plotted in Fig. 6.
Figure 6: Evolution of deceleration parameter $q$ with redshift $z$ for the WSQ model and $\Lambda$CDM, plotted with best-fit parameters obtained for the SNIa+OHD datasets.
@@
$r=\frac{1}{a}\frac{d^{3}a}{dt^{3}}\left[\frac{1}{a}\frac{da}{dt}\right]^{-3}=\frac{\dddot{a}}{aH^{3}}.$ (4.4)
We can represent $r$ in terms of the dimensionless Hubble parameter $h$ while introducing a new variable $N=\text{ln}(a)$, we obtain,
$r=\frac{1}{2h^{2}}\frac{d^{2}h^{2}}{dN^{2}}+\frac{3}{2h^{2}}\frac{dh^{2}}{dN}+1.$ (4.5)
Figure 7: The statefinder trajectory is plotted using the best-fit parameters from the SNIa + OHD datasets. The fixed point of $\Lambda$CDM is also marked.
@@
$s=\frac{r-1}{3(q-1/2)}.$ (4.6)
On expressing $q$ presented in Eq. (4.2) in terms of $N$ and substituting in Eq. (4.6), we obtain
$s=-\frac{\frac{1}{2h^{2}}\frac{d^{2}h^{2}}{dN^{2}}+\frac{3}{2h^{2}}\frac{dh^{2}}{dN}}{\frac{3}{2h^{2}}\frac{dh^{2}}{dN}+\frac{9}{2}}.$ (4.7)
@@
The statefinder trajectory distinguishes the behaviour of different types of dark energy models. The trajectory is plotted in Fig. 7. The statefinder trajectory indicates the quintessence behavior of the field. The present value of the statefinder diagnostic pair is obtained as ($r,s$) = ($0.8309,0.0562$), different from the $\Lambda$CDM fixed point ($1,0$).
We now study the physical properties of the scalar field, i.e., how the scalar field and its potential evolves with redshift. We present the qualitative evolution of the potential $V(\phi)$ as a function of redshift in Fig. 8. The potential is higher at earlier times as the scalar field rolls down the potential. The qualitative evolution of the scalar field can be studied from the quantity $\Delta\phi(z)=\phi(z)-\phi_{0}$ obtained from Eq. (2.8) and using the best fit parameters. This is presented in Fig. 9. The scalar field was initially lower than the present value and then increased at later times.
Figure 8: The evolution of the scalar field potential $V(\phi)$ plotted using the best-fit parameters from the SNIa + OHD dataset.
Figure 9: The qualitative evolution of the quantity $\Delta\phi=\phi-\phi_{0}$ plotted using the best-fit parameters from the SNIa + OHD dataset.
5 Dynamical System Analysis
In order to gain an understanding of the global behaviour any cosmological model, dynamical system analysis is particularly useful [80, 81], especially for studying its asymptotic behaviour. This requires a careful choice of dynamical variables to study the phase-space dynamics of the system.
$x=\frac{\dot{\phi}\kappa}{\sqrt{6}H},\hskip 14.45377pty=\frac{\kappa\sqrt{V(\phi)}}{\sqrt{3}H}.$ (5.1)
The dark energy or scalar field density is then defined as,
$\Omega_{\phi}=\frac{\kappa^{2}\rho_{\phi}}{3H^{2}}=x^{2}+y^{2}.$ (5.2)
The effective equation of state of the background cosmological fluid would be,
$\omega_{eff}=\frac{P_{\phi}+P_{m}}{\rho_{\phi}+\rho_{m}}=x^{2}-y^{2}$ (5.3)
From Eq. (2.6) and (2.7), we get the relation,
$\frac{\dot{H}}{H^{2}}=-\frac{3}{2}(1+x^{2}-y^{2}).$ (5.4)
Using this equation and Eq. (5.1), we differentiate these variables with respect to $N=\text{ln}(a)$, to obtain a system of autonomous differential equations. The dynamical variables $x$ and $y$ cannot close the system of autonomous equations without considering another dynamical variable $\lambda$, which is defined as,
$\lambda=-\frac{V_{,\phi}}{\kappa V(\phi)}=\left(1-\frac{V(\phi)}{2V_{0}}\right).$ (5.5)
The system of autonomous differential equations are then given by,
$\frac{dx}{dN}=-3x+\frac{\sqrt{6}}{2}\lambda y^{2}+\frac{3}{2}x(1+x^{2}-y^{2})$ (5.6)
$\frac{dy}{dN}=-\frac{\sqrt{6}}{2}\lambda xy+\frac{3}{2}y(1+x^{2}-y^{2})$ (5.7)
$\frac{d\lambda}{dN}=\sqrt{6}x\lambda(1-\lambda).$ (5.8)
Before studying the dynamics, we need to make sure that the phase space of this system is compact. The variables $x$ and $y$ are constraint by Friedmann equation (2.6). The variable $\lambda$ is also compact, from the fact that the minimum value of $V(\phi)$ is zero and the maximum value is effectively $2V_{0}$. Thus the phase space of this system is represented by the positive-$y$ half cylinder from $\lambda=0$ to $\lambda=1$. The phase space portrait for this system is given in Fig. 10. The fixed points of the system can be obtained by setting Eq. (5.6), (5.7) and (5.8) to zero. These critical points and their properties are outlined in Table 3.
Refer to caption Refer to caption
Figure 10: The phase space trajectory for the WSQ model, depicting a single streamline on the left and multiple streamlines converging to the point $C$ on the right.
Point $x_{c}$ $y_{c}$ $\lambda_{c}$ Accel. $\Omega_{\phi}$ $\omega_{eff}$ $\omega_{\phi}$
$O_{\lambda}$ 0 0 Any No 0 0 -
$A_{\pm}$$\pm 1$ 0 1 No 1 1 1
$B$ 0 1 0 Yes 1 -1 -1
$C$ $\frac{1}{\sqrt{6}}$ $\sqrt{\frac{5}{6}}$ 1 Yes 1 -$\frac{2}{3}$$\frac{2}{3}$
Table 3: Critical Points
It is to be noted that the dynamical variables $x$, $y$ and $\lambda$ are constrained to lie in the ranges $-1\leq x\leq 1$, $0\leq y\leq 1$, and $0\leq\lambda\leq 1$ as per their definition (Eq. 5.1) and Eq. (2.6). So critical points that lie outside this range have not been considered. Additionally, when $y\rightarrow 0$, $\lambda\rightarrow 1$ according to Eq. (5.5), and critical points which do not satisfy this constraint have also been not considered. In order to study the stability of these fixed points $(x,y,\lambda)=(x_{c},y_{c},\lambda_{c})$, we consider linear perturbations about these points $\delta x,\delta y$ and $\delta\lambda$ such that they satisfy,
$\frac{d}{dN}\begin{pmatrix}\delta x\\ \delta y\\ \delta y\end{pmatrix}=J\begin{pmatrix}\delta x\\ \delta y\\ \delta y\end{pmatrix}.$ (5.9)
where the Jacobian matrix $J$ is given by,
$J=\begin{pmatrix}\frac{\partial f}{\partial x}&\frac{\partial f}{\partial y}&\frac{\partial f}{\partial\lambda}\\[9.0pt] \frac{\partial g}{\partial x}&\frac{\partial g}{\partial y}&\frac{\partial g}{\partial\lambda}\\[9.0pt] \frac{\partial h}{\partial x}&\frac{\partial h}{\partial y}&\frac{\partial h}{\partial\lambda}\end{pmatrix}_{x=x_{c},y=y_{c},\lambda=\lambda_{c}},$ (5.10)
where $f(x,y,\lambda)$, $g(x,y,\lambda)$ and $h(x,y,\lambda)$ are the RHS of Eq. (5.6),(5.7),(5.8) respectively. The stability of the critical points is determined by the eigenvalues of this Jacobian matrix at these critical points. If all eigenvalues are negative, the critical point is a future attractor, which is stable. If all the eigenvalues are positive, the corresponding point is a past attractor and is unstable and if some eigenvalues are positive and some negative, the critical point is a saddle point. If the eigenvalues are complex, the critical points would be stable spiral or unstable spiral depending on the real parts of the eigenvalues. These points with eigenvalues having non-zero real part are called $hyperbolic$ points [82]. The linear stability theory fails if any of the eigenvalues is zero and other methods have to be employed to determine the stability of the system at this $non-hyperbolic$ critical point. The eigenvalues of the critical points presented in Table 3 and their stability are given in Table 4,
Point Eigenvalues Nature of critical point Stability
$O_{\lambda}$ $\left(-\frac{3}{2},\frac{3}{2},0\right)$ Non-hyperbolic Saddle
$A_{\pm}$ $\left(3,\pm\sqrt{6},\frac{1}{2}(6\pm\sqrt{6})\right)$ Hyperbolic Unstable/Saddle
$B$ $\left(-3,-3,0\right)$ Non-hyperbolic Saddle
$C$ $\left(-\frac{5}{2},-2,-1\right)$ Hyperbolic Stable
Table 4: Critical Points and their eigenvalues
For non-hyperbolic critical points, there are usually two approaches to study their stability. One is using Lyapunov functions[83, 84], and the other is the Centre Manifold theorem. Here we have employed the centre manifold theorem to analyze the stability of the non-hyperbolic point $B$.
5.1 Analysis of critical points
  • Points $O_{\lambda}$ : This set of points correspond to the $\lambda$-axis, i.e., the $\lambda$-axis acts as a critical line. On this line, the Universe is matter-dominated, with the scalar field energy density (and pressure) vanishing at this critical line, leaving $\omega_{\phi}$ undetermined and the effective equation of state (EoS) $\omega_{eff}$ equal to the matter equation of state, i.e., zero. This gives a non-accelerating solution because of the absence of the scalar field. From the eigenvalues of the critical points, one eigenvalue vanishes and hence this set of points is non-hyperbolic and linear stability is inapplicable. But from the remaining two eigenvalues, one is positive and the other is negative. Thus, the critical points $O_{\lambda}$ is asymptotically unstable and thus will act as a saddle line.
  • Points $A_{\pm}$ : These two points (±$(\pm 1,0,1)$correspond to scalar field dominated solutions where the kinetic component of the scalar field dominates. The Universe is completely dominated by the scalar field and results in an effective EoS $\omega_{eff}=1$, which characterizes a stiff fluid dominated solution. Both points are associated with non-accelerated expansion of the Universe. The eigenvalues corresponding to these points suggest that they are unstable, with $A_{+}$ being an unstable point (a past attractor) and $A_{-}$ being a saddle point.
  • Point $B$ : This point $(0,1,0)$ corresponds to a cosmological constant-like dominated solution where the Universe undergoes a de Sitter accelerated expansion, with the effective EoS $\omega_{eff}=-1$ and the scalar field completely dominating the Universe. This point represents the scalar field potential energy dominated solution with vanishing kinetic energy. From Table 4, it is evident that this point is non-hyperbolic, because of the vanishing eigenvalue. Hence, linear stability is not applicable and to determine whether this point could act as a stable future attractor, we apply Center manifold theorem. Through this approach (Sec. 5.2), we find that this critical point is asymptotically unstable, indicating that it functions as a saddle point in the dynamics of the system.
  • Point $C$ : This point $\left(\frac{1}{\sqrt{6}},\sqrt{\frac{5}{6}},1\right)$ also represents a solution where the Universe is entirely dominated by the scalar field as the scalar field energy density $\Omega_{\phi}=1$. Unlike Point $B$, this point has non-zero contributions from both the kinetic and potential energy components of the scalar field. This solution gives an effective EoS $\omega_{eff}=-\frac{2}{3}$, resulting in accelerated expansion of the Universe. Since the eigenvalues are non-zero, linear stability theory can be applied to this point. With all three eigenvalues being negative, this indicates that the point is a stable future attractor.
5.2 Centre Manifold Approach to the Point $B$
The critical point $B$ $(0,1,0)$ has eigenvalues $(-3,-3,0)$ and linear stability theorem cannot be applied to study the stability of this point. In order to apply the centre manifold theorem, we first shift the critical point to the origin, by transforming the dynamical variables in Eq. (5.1) into a new set of variables,
$X=x,\hskip 14.45377ptY=y-1,\hskip 14.45377ptZ=\lambda.$ (5.11)
Then the derivatives with respect to $N=\text{ln}(a)$ for these new variables are given by,
$\frac{dX}{dN}=-3X+\frac{\sqrt{6}}{2}Z(Y+1)^{2}+\frac{3}{2}X(1+X^{2}-(Y+1)^{2})$ (5.12)
$\frac{dY}{dN}=-\frac{\sqrt{6}}{2}ZX(Y+1)+\frac{3}{2}(Y+1)(1+X^{2}+(Y+1)^{2})$ (5.13)
$\frac{dZ}{dN}=\sqrt{6}XZ(1-Z).$ (5.14)
Now we find the Jacobian for this set of autonomous equations at the origin (0,0,0). We obtain the Jacobian as
$J\big{|}_{X=0,Y=0,Z=0}=\begin{pmatrix}-3&0&\frac{\sqrt{6}}{2}\\ 0&-3&0\\ 0&0&0\end{pmatrix}.$ (5.15)
Now, we introduce a new set of variables such that,
$\begin{pmatrix}U\\ V\\ W\end{pmatrix}=S^{-1}\begin{pmatrix}X\\ Y\\ Z\end{pmatrix},$ (5.16)
where $S$ is the matrix formed by the eigenvectors of the Jacobian matrix $J$. We obtain the matrix $S$ as,
$S=\begin{pmatrix}1&0&\frac{1}{\sqrt{6}}\\ 0&1&0\\ 0&0&1\end{pmatrix}$ (5.17)
Then, from Eq. (5.16), we get the new set of variables as,
$U=X-\frac{Z}{\sqrt{6}},\hskip 14.45377ptV=Y,\hskip 14.45377ptW=Z.$ (5.18)
Under this transformation, Eq.(5.12),(5.13) and (5.14)can be rewritten as,
$\begin{split}\frac{dU}{dN}=&-3\left(U+\frac{W}{\sqrt{6}}\right)+\frac{\sqrt{6}}{2}W(V+1)^{2}\\ &+\frac{3}{2}\left(U+\frac{W}{\sqrt{6}}\right)\left[1-\left(U+\frac{W}{\sqrt{6}}\right)^{2}+(V+1)^{2}\right]-\left(U+\frac{W}{\sqrt{6}}\right)W(1-W),\end{split}$ (5.19)
$\frac{dV}{dN}=$ $-\frac{\sqrt{6}}{2}W\left(U+\frac{W}{\sqrt{6}}\right)(V+1)+\frac{3}{2}(V+1)\left[1-\left(U+\frac{W}{\sqrt{6}}\right)^{2}+(V+1)^{2}\right],$ (5.20)
$\frac{dW}{dN}=$ $\sqrt{6}\left(U+\frac{W}{\sqrt{6}}\right)W(1-W).$ (5.21)
According to the centre manifold theory, these equations can be represented in the form,
$\frac{dU}{dN}=AU+F(U,V,W)$ (5.22)
$\frac{dV}{dN}=BV+G(U,V,W)$ (5.23)
$\frac{dW}{dN}=CW+H(U,V,W)$ (5.24)
where we obtain $A=-3$, $B=-3$ and $C=0$. Then, by centre manifold theorem, there exists a centre manifold such that the dynamics of the system reduced to the centre manifold is given by,
$\frac{dW}{dN}=CW+H(h_{1}(W),h_{2}(W),W),$ (5.25)
for sufficiently small $W$. Here we have assumed $U=h_{1}(W)$ and $V=h_{2}(W)$. This should satisfy the quasilinear partial differential equations,
$\frac{dh_{1}(W)}{W}\left(CW+H(h_{1}(W),h_{2}(W),W)\right)-\left(Ah_{1}(W)+F(h_{1}(W),h_{2}(W),W)\right)=0$ (5.26)
$\frac{dh_{2}(W)}{W}\left(CW+H(h_{1}(W),h_{2}(W),W)\right)-\left(Bh_{2}(W)+G(h_{1}(W),h_{2}(W),W)\right)=0$ (5.27)
Since an exact knowledge of the centre manifold is not necessary, it suffices to approximate the centre manifold[85] and hence it is customary to assume the expansions,
$h_{1}(W)=a_{2}W^{2}+a_{3}W^{3}+\mathcal{O}(W^{4})$ (5.28)
$h_{2}(W)=b_{2}W^{2}+b_{3}W^{3}+\mathcal{O}(W^{4}).$ (5.29)
Substituting this into Eq. (5.26) and (5.27), we can equate the coefficients of each power of $W$ equal to zero to obtain the values of the coefficients $a_{2}$, $a_{3}$, $b_{2}$ and $b_{3}$. We obtain $a_{2}=-\dfrac{1}{3\sqrt{6}}$, $a_{3}=\dfrac{2}{3\sqrt{6}}$, $b_{2}=-\dfrac{1}{12}$ and $b_{3}=\dfrac{1}{18}$. Thus Eq. (5.25) becomes,
$\begin{split}\frac{dW}{dN}&=W^{2}-W^{3}+\sqrt{6}\left(a_{2}W^{2}+a_{3}W^{3}\right)W-\sqrt{6}\left(a_{2}W^{2}+a_{3}W^{3}\right)W^{2}\\ &=W^{2}-\frac{4}{3}W^{3}+W^{4}+\mathcal{O}(W^{5})\end{split}$ (5.30)
From Eq. (5.30), it can be concluded that for sufficiently small W, the system is unstable as the lowest order term has even parity. Hence the critical point $(0,1,0)$ is unstable and acts as a saddle point.
To conclude the results of the dynamical system analysis, we obtained one stable critical point (Point $C$), two saddle points and a saddle line, and an unstable critical point. Thus, the Universe is asymptotically stable within the WSQ model and attains an accelerated expansion. Since the WSQ model fits observation and predicts parameter values consistent with that of the standard model, and is dynamically stable, this model may be considered as a potential alternative to the standard model.
@@ 6 Conclusion
In the presented work, we examined the scalar field model with Woods-Saxon like potential. In an earlier work, Oikonomou and Chatzarakis demonstrated that early inflation can be successfully realized using a scalar field model with a similar potential. In this work, we study the implications of such a scalar field in the late Universe in producing the observed acceleration.
We tested the model with the latest observational data using the data combination SNIa + OHD and adopting Bayesian Inference. Our analysis shows that the model is consistent with the observational data and the obtained values of parameters are also consistent with that obtained with $\Lambda$CDM for the same dataset. The obtained best-fit values of model parameters are $H_{0}=68.10\pm 0.74$, $\Omega_{m_{0}}=0.31\pm 0.01$ and $y_{0}=0.83\pm 0.01$, which is in close agreement with $\Lambda$CDM predictions. The WSQ model is preferred over the $\Lambda$CDM according to $\chi^{2}_{min}$ and AIC, while according to the BIC criteria, the $\Lambda$CDM is slightly preferred over the WSQ model, because of the extra parameter.
The study of evolution of different cosmological parameters also revealed the consistency of the WSQ model in explaining the late-time background evolution of the Universe. The evolution of $\omega$ shows a freezing model of quintessence. Age of the Universe is estimated to be $13.6$ Gyr, which is slightly less than that predicted by the $\Lambda$CDM. The present value of the deceleration parameter obtained with the WSQ model is negative, showing that the model predicts present day accelerated expansion. The Universe transitioned from decelerating to accelerating is at a redshift $z_{T}=0.662$. The statefinder trajectory reveals the quintessence behaviour of the field.
We performed dynamical system analysis to test the stability of the Universe in the WSW model. Studying the evolution of the dynamical variables $\kappa\dot{\phi}/\sqrt{6}H$, $\kappa\sqrt{V(\phi)}/\sqrt{3}H$ and $\lambda=(1-V(\phi)/2V_{0})$ from their autonomous coupled differential equations, we obtained multiple critical points, of which only one was a stable attractor according to linear stability theorem. One non-hyperbolic critical point was studied using the centre manifold approach and was found to be unstable. In conclusion, the WSQ model does allow for a future attractor and the Universe is stable within the WSQ model.
In summary, the model is consistent with the observations to good extent and is dynamically stable. Thus, this model can be considered as a potential alternative to the $\Lambda$CDM to explain the late phase acceleration of the Universe. This complements the findings of [69], which studied the early phase inflation and reheating with a scalar field model with a similar Woods-Saxon potential, reinforcing the possibility of such a component driving accelerated expansion in the late Universe.
@@
Acknowledgments
The authors of the manuscript are thankful to the Indian Institute of Technology Kanpur for providing the Param Sanganak high-performance computation facility for faster execution of the Python program. One of the authors, Sarath Nelleri, is thankful to the Indian Institute of Technology Kanpur for providing the Institute Postdoctoral Fellowship. The author Sreerag Radhakrishnan is thankful to Mrs. Sanitha A for fruitful discussions on the topic.